### v1

# Finite groups with some subgroups of prime power order satisfying the partial $\Pi$-property

Abstract
Let $H$ be a subgroup of a finite group $G$. We say that $H$ satisfies the partial $\Pi$-property in $G$ if there exists a $G$-chief series $\varGamma_{G}:1=G_{0}<G_{1}<\cdot\cdot\cdot<G_{n}=G$ of $G$ such that $|G/G_{i-1}:N_{G/G_{i-1}}(HG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})|$ is a $\pi(HG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})$-number for every $G$-chief factor $G_{i}/G_{i-1}$ of $\varGamma_{G}$, $1\leq i\leq n$. In this paper, we investigate the structure of a finite group $G$ under the assumption that some subgroups of prime power order satisfy the partial $\Pi$-property.
Keywords: Finite group, $p$-soluble group, partial $\Pi$-property.
Mathematics Subject Classification (2020): 20D10, 20D20.
1 Introduction
All groups considered in this paper are finite.
The present paper is a further contribution to the research project that analyses the impact of the embedding of some relevant families of subgroups on the structure of a group. The subgroup embedding property studied here is the partial $\Pi$-property introduced by Chen and Guo in [4], which generalises a large number of known embedding properties (see [4, Section 7]), and has raised a considerable interest in recent years.
Definition 1.1.
We say that a subgroup $H$ of a group $G$ satisfies the partial $\Pi$-property in $G$ if there exists a $G$-chief series $\varGamma_{G}:1=G_{0}<G_{1}<\cdot\cdot\cdot<G_{n}=G$ of $G$ such that $|G/G_{i-1}:N_{G/G_{i-1}}(HG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})|$ is a $\pi(HG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})$-number for every $G$-chief factor $G_{i}/G_{i-1}$ of $\varGamma_{G}$, $1\leq i\leq n$.
As usual, $\pi(G)$ denotes the set of all primes dividing the order of a group $G$.
Our aim here is to solve three problems concerning the structural influence of the partial $\Pi$-property of some subgroups of prime power order and confirm that the partial $\Pi$-property is a true generalisation of the partial CAP-property studied in [1, 3, 2, 8].
Chen and Guo proved in [4, Proposition 1.6] that a group $G$ in which every subgroup of order $p$, $p$ a prime, and every cyclic subgroup of order $4$ (if $p=2$ and $P$ is not quaternion-free), satisfy the partial $\Pi$-property in $G$ is $p$-supersoluble, that is, every chief factor of $G$ divisible by $p$ is cyclic.
Our first main result provides a structural information of a group $G$ in which every subgroup of order $p^{2}$ satisfies the partial $\Pi$-property. Without loss of generality, we may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
Theorem A.
Le $p$ be a prime, and let $P$ be a Sylow $p$-subgroup of a group $G$. Assume that $O_{p^{\prime}}(G)=1$ and $|P|\geq p^{2}$. If every subgroup of $P$ of order $p^{2}$ satisfies the partial $\Pi$-property in $G$, then $P=O_{p}(G)$ and $G=P\rtimes H$, the semidirect product of $P$ and $H$, where $H\in{\rm Hall}_{p^{\prime}}(G)$, and one of the following statements holds:
(1) $G$ is $p$-supersoluble;
(2) $P$ is a minimal normal subgroup of $G$ of order $p^{2}$;
(3) $|P|\geq p^{4}$ and $H$ is cyclic. Furthermore, $P=V_{1}\times\cdot\cdot\cdot\times V_{s}$, $s\geq 2$, where $V_{i}$ is an $H$-isomorphic irreducible $\mathbb{F}_{p}[H]$-module of dimension $2$ for all $1\leq i\leq s$, where $\mathbb{F}_{p}$ is the finite field of $p$-elements.
Chen and Guo proved in [4, Proposition 1.4] that if the maximal subgroups of the Sylow $p$-subgroups of a group $G$ satisfy the $\Pi$-property in $G$, then either $G$ is $p$-supersoluble or the order of the Sylow $p$-subgroups of $G$ is at most $p$.
Our second main result describes the structure of a group $G$ in which the second maximal subgroups of the Sylow $p$-subgroups have the partial $\Pi$-property in $G$. We may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
Theorem B.
Let $P$ be a Sylow $p$-subgroup of a group $G$. Assume that $O_{p^{\prime}}(G)=1$ and $|P|\geq p^{2}$. If every $2$-maximal subgroup of $P$ satisfies the partial $\Pi$-property in $G$, then one of the following statements holds:
(1) $G$ is $p$-supersoluble;
(2) If $G$ is $p$-soluble and $|P|=p^{2}$, then $P=O_{p}(G)$ is a minimal normal subgroup of $G$;
(3) $p=2$ and $P$ is isomorphic to $Q_{8}$;
(4) $|P|\geq p^{3}$ and $G=P\rtimes H$, where $H\in{\rm Hall}_{p^{\prime}}(G)$ is cyclic. Furthermore, $\Phi(P)$ is the intersection of all $2$-maximal subgroups of $P$, and $P/\Phi(P)=V_{1}\times\cdot\cdot\cdot\times V_{s}$, where $V_{i}$ is an $H$-isomorphic irreducible $\mathbb{F}_{p}[H]$-module of dimension $2$ for all $1\leq i\leq s$.
A remarkable extension of Chen and Guo’s results is due to the first author, Liu and Chen.
Theorem 1.2 ([9, Theorem 1.3]).
Let $P$ be a Sylow $p$-subgroup of a group $G$, and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$, and every cyclic subgroup of $P$ of order $4$ (when $d=2$ and $P$ is not quaternion-free) satisfy the partial $\Pi$-property in $G$. Then $G$ is $p$-soluble of $p$-length at most $1$.
Our third main result describes the structure of the groups of $p$-rank greater than $1$ satisfying the hypotheses of Theorem 1.2. Recall that the $p$-rank of a $p$-soluble group $G$ is the largest integer $k$ such that $G$ has a chief factor of order $p^{k}$. Note that a $p$-soluble group $G$ is $p$-supersoluble if and only if the $p$-rank of $G$ is equal to $1$.
Theorem C.
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$, and assume further that every cyclic subgroup of $P$ of order $4$ satisfies the partial $\Pi$-property in $G$ when $d=2$ and $P$ is not quaternion-free. If the $p$-rank of $G$ is greater than $1$, then $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Furthermore, if $V$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P/\Phi(P)$, and write
$dim_{\mathbb{F}_{p}}V=k$, $n=\log_{p}d-\log_{p}|\Phi(P)|$, and $m=\log_{p}|P/\Phi(P)|$,
then the following statements hold:
(1) $P/\Phi(P)$ is a homogeneous $\mathbb{F}_{p}[H]$-module and every irreducible submodule of $P/\Phi(P)$ is not absolutely irreducible;
(2) $n\geq k\geq 2$, $k$ divides $\gcd(m,n)$;
(3) $H$ is cyclic.
2 Preliminaries
Before we can begin our proofs of Theorems A,  B and C, we need a number of preliminary results. The first of these is very useful in induction arguments.
Lemma 2.1 ([4, Lemma 2.1]).
Let $H$ be a subgroup of a group $G$ and $N\unlhd G$. If either $N\leq H$ or $\gcd(|H|,|N|)=1$ and $H$ satisfies the partial $\Pi$-property in $G$, then $HN/N$ satisfies the partial $\Pi$-property in $G/N$.
The next lemma is crucial to prove useful results on the partial $\Pi$-property.
Lemma 2.2 ([10, Lemma 2.3]).
Let $H$ be a $p$-subgroup of a group $G$ and $N$ be a normal subgroup of $G$ containing $H$. If $H$ satisfies the partial $\Pi$-property in $G$, then $G$ has a chief series
$\varOmega_{G}:1=G^{*}_{0}<G^{*}_{1}<\cdot\cdot\cdot<G^{*}_{r}=N<\cdot\cdot\cdot<G^{*}_{n}=G$
passing through $N$ such that $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ $(1\leq i\leq n)$ of $\varOmega_{G}$.
The order of the minimal normal subgroups of groups with subgroups of prime power order satisfying the partial $\Pi$-property is quite constrained.
Lemma 2.3 ([9, Lemma 2.1]).
Let $P$ be a Sylow $p$-subgroup of a group $G$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$. Then:
(1) Every minimal normal subgroup of $G$ is either a $p^{\prime}$-group or a $p$-group of order at most $d$.
(2) If $G$ has a minimal normal subgroup of order $d$, then every minimal normal $p$-subgroup of $G$ has order $d$.
Recall that the supersoluble hypercenter $Z_{{\mathcal{U}}}(G)$ of a group $G$ is the product of all normal subgroups $H$ of $G$, such that all $G$-chief factors under $H$ have prime order, and the $p$-supersoluble hypercenter $Z_{{\mathcal{U}}_{p}}(G)$ is the product of all normal subgroups $H$ of $G$, such that all $p$-$G$-chief factors under $H$ have order $p$ for some fixed prime $p$. Note that a normal $p$-subgroup is contained in $Z_{{\mathcal{U}}}(G)$ if and only if it is contained in $Z_{{\mathcal{U}}_{p}}(G)$.
Lemma 2.4 ([9, Lemma 2.11]).
Let $P$ be a normal $p$-subgroup of a group $G$. If all cyclic subgroups of $P$ of order $p$ or $4$ (when $P$ is not quaternion-free) are contained in $Z_{{\mathcal{U}}}(G)$, then $P\leq Z_{{\mathcal{U}}}(G)$.
If $P$ is either an odd order $p$-group or a quaternion-free $2$-group, then we use $\Omega(P)$ to denote the subgroup $\Omega_{1}(P)$. Otherwise, $\Omega(P)=\Omega_{2}(P)$.
Our next lemma is a consequence of [5, Lemma 2.8].
Lemma 2.5.
Let $P$ be a normal $p$-subgroup of a group $G$ and $D$ a Thompson critical subgroup of $P$ (see [7, page 186]). If $P/\Phi(P)\leq Z_{{\mathcal{U}}}(G/\Phi(P))$ or $\Omega(D)\leq Z_{{\mathcal{U}}}(G)$, then $P\leq Z_{{\mathcal{U}}}(G)$.
Lemma 2.6 ([5, Lemma 2.10]).
Let $D$ be a Thompson critical subgroup of a non-trivial $p$-group $P$.
(1) If $p>2$, then the exponent of $\Omega_{1}(D)$ is $p$.
(2) If $P$ is an abelian $2$-group, then the exponent of $\Omega_{1}(D)$ is $2$.
(3) If $p=2$, then the exponent of $\Omega_{2}(D)$ is at most $4$.
Lemma 2.7 ([12, Lemma 3.1]).
Let $P$ be a non-abelian quaternion-free $2$-group. Then $P$ has a characteristic subgroup of index $2$.
Lemma 2.8 ([11, Lemma 2.10]).
Let $p$ be a prime, $E$ a normal subgroup of a group $G$ such that $p$ divides the order of $E$. Then $E\leq Z_{{\mathcal{U}}_{p}}(G)$ if and only if $E/\Phi(E)\leq Z_{{\mathcal{U}}_{p}}(G/\Phi(E))$.
Lemma 2.9.
Let $p$ be a prime. Assume that $N$ is a minimal normal subgroup of a group $G$ and $K$ a subgroup of $G$. Assume that $|N|=|K|=p$. If $NK$ satisfies the partial $\Pi$-property in $G$, then $K$ also satisfies the partial $\Pi$-property in $G$.
Proof.
If $K=N$, there is nothing to prove. Hence we can assume that $NK$ has order $p^{2}$. There exists a chief series $\varGamma_{G}:1=G_{0}<G_{1}<\cdot\cdot\cdot<G_{n}=G$ of $G$ such that $|G/G_{i-1}:N_{G/G_{i-1}}(NKG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})|$ is a $p$-number, for every $G$-chief factor $G_{i}/G_{i-1}$ of $\varGamma_{G}$, $1\leq i\leq n$ Our goal is to show that if $G_{i}/G_{i-1}$ is a chief factor of $\varGamma_{G}$, it follows that $|G/G_{i-1}:N_{G/G_{i-1}}(KG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})|$ is a $p$-number. Obviously we may assume that $|(KG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|\neq 1$.
Assume that $|(NKG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|=p^{2}$, then $NK\cap G_{i-1}=1$ and $NKG_{i-1}\leq G_{i}$. Note that $G_{i}=NG_{i-1}$. Hence $|(NKG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|=|(NKG_{i-1}/G_{i-1})\cap(NG_{i-1}/G_{i-1})|=|NG_{i-1}/G_{i-1}|=p$, contrary to our assumption. Hence $|(NKG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|\leq p$ and so $|(KG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|=p=|(NKG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})|$. Thus $(KG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})=(NKG_{i-1}/G_{i-1})\cap(G_{i}/G_{i-1})$ and therefore $|G/G_{i-1}:N_{G/G_{i-1}}(KG_{i-1}/G_{i-1}\cap G_{i}/G_{i-1})|$ is a $p$-number. Consequently, $K$ satisfies the partial $\Pi$-property in $G$. ∎
Lemma 2.10.
Let $P\in{\rm Syl}_{p}(G)$, $d$ be a power of $p$ with $p\leq d\leq|P|$, and let $N$ be a minimal normal subgroup of $G$ whose order is divisible by $d$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$. Then $N$ is an elementary abelian subgroup of $G$ of order $d$.
Proof.
Let $H$ be a normal subgroup of $P$ of order $d$ such that $H\leq N$. By hypothesis, $H$ satisfies the partial $\Pi$-property in $G$. It follows from Lemma 2.2 that $|G:N_{G}(H)|$ is a $p$-number, and so $H\unlhd G$. The minimality of $N$ yields that $H=N$. Thus $N$ is an elementary abelian subgroup of $G$ of order $d$. ∎
Lemma 2.11.
Let $P$ be a normal Sylow $p$-subgroup of a group $G$. Suppose that a $2$-maximal subgroup $H$ of $P$ satisfies the partial $\Pi$-property in $G$. Then $H$ is a partial CAP-subgroup of $G$.
Proof.
Since $H\leq P\unlhd G$, it follows from Lemma 2.2 that $G$ has a chief series
$\varOmega_{G}:1=G^{*}_{0}<G^{*}_{1}<\cdots<G^{*}_{r}=P<\cdots<G^{*}_{n}=G$
passing through $P$ such that $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. Observe that $Z(P)\cap G^{*}_{1}\neq 1$. Since $Z(P)$ is normal in $G$ and $G^{*}_{1}$ is a minimal normal subgroup of $G$, it follows that $G^{*}_{1}\leq Z(P)$. In particular, $H\cap G^{*}_{1}$ is normalised by $P$. Since $|G:N_{G}(H\cap G^{*}_{1})|$ is a $p$-number, it follows that $H\cap G^{*}_{1}$ is normal in $G$. Therefore either $H\cap G^{*}_{1}=1$ or $H\cap G^{*}_{1}=G^{*}_{1}$.
Assume that $H\cap G^{*}_{1}=1$. For every $G$-chief factor $G_{i}/G_{i-1}$ with $1<i\leq r$, we see that either $P=HG^{*}_{i-1}$ or $HG^{*}_{i-1}$ is a maximal subgroup of $P$. In both cases, $P$ normalises $HG^{*}_{i-1}$ and hence $P\leq N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})$. Since $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number, it follows that $HG^{*}_{i-1}\cap G^{*}_{i}$ is a normal subgroup of $G$. Therefore $H$ covers or avoids $G^{*}_{i}/G^{*}_{i-1}$. This implies that $H$ is a partial CAP-subgroup of $G$, as desired. Hence we can assume that $H\cap G^{*}_{1}=G^{*}_{1}$. By Lemma 2.1, the hypotheses are inherited by $G/G^{*}_{1}$. By induction, $H/G^{*}_{1}$ is is a partial CAP-subgroup of $G/G^{*}_{1}$. Hence $G/G^{*}_{1}$ has a chief series
$\varGamma_{G/G^{*}_{1}}:1=G_{1}/G^{*}_{1}<G_{2}/G^{*}_{1}<\cdots<G_{n}/G^{*}_{1}=G/G^{*}_{1}$
such that $H/G^{*}_{1}$ covers or avoids every $G/G^{*}_{1}$-chief factor $G_{j}/G^{*}_{1}\big{/}G_{j-1}/G^{*}_{1}$ of $\varGamma_{G/G^{*}_{1}}$, $2\leq j\leq n$. Therefore, $G_{j}/G^{*}_{1}\leq HG_{j-1}/G^{*}_{1}$ or $(H\cap G_{j})/G^{*}_{1}\leq G_{j-1}/G^{*}_{1}$. Consequently, $G_{j}\leq HG_{j-1}$ or $H\cap G_{j}\leq G_{j-1}$. This means that $H$ covers or avoids every $G$-chief factor of a chief series of $G$. Thus $H$ is a partial CAP-subgroup of $G$, as desired. ∎
Recall that a subgroup $H$ of $G$ is said to be complemented in a group $G$ if there exists a subgroup $K$ of $G$ such that $G=HK$ and $H\cap K=1$. In this case, $K$ is called a complement of $H$ in $G$.
Lemma 2.12.
Let $G$ be a group with an elementary abelian normal Sylow $p$-subgroup $P$. Suppose that $H$ is a subgroup of $P$. Then the following statements are equivalent:
(1) $H$ satisfies the partial $\Pi$-property in $G$.
(2) $H$ is complemented in $G$.
Proof.
(1$\Rightarrow$(2). We argue by induction on $|G|$. If $H=1$ there is nothing to prove. By the Schur-Zassenhaus Theorem ([6, Theorem A.11.3]), $P$ is complemented in $G$ by a Hall $p^{\prime}$-subgroup $U$ of $G$. Hence we may assume that $H<P$. By Lemma 2.2, $G$ has a chief series
$\varOmega_{G}:1=G^{*}_{0}<G^{*}_{1}<\cdot\cdot\cdot<G^{*}_{r}=P<\cdot\cdot\cdot<G^{*}_{n}=G$
passing through $P$ such that $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. In particular, $|G:N_{G}(H\cap G^{*}_{1})|$ is a $p$-number, and so $G=PN_{G}(H\cap G^{*}_{1})$. Since $P$ is elementary abelian, we have $(H\cap G_{1}^{*})^{G}=H^{P}\cap G_{1}^{*}=H\cap G_{1}^{*}$. The minimality of $G_{1}^{*}$ implies that $H\cap G_{1}^{*}=G_{1}^{*}$ or $1$. If $H\cap G_{1}^{*}=G_{1}^{*}$, then $G_{1}^{*}\leq H$. By Lemma 2.1, $H/G_{1}^{*}$ satisfies the partial $\Pi$-property in $G/G_{1}^{*}$. By induction, $H/G_{1}^{*}$ has a complement in $G/G_{1}^{*}$, $K/G_{1}^{*}$ say. Thus $G=HK$ and $H\cap K=G_{1}^{*}$. Hence $P=H(P\cap K)$ and $P\cap K\unlhd G$. By Maschke’s theorem ([6, Theorem A.11.4]), there exists a normal subgroup $L$ of $G$ such that $P\cap K=G_{1}^{*}\times L$. In this case, $LU$ is a complement of $H$ in $G$, as required.
If $H\cap G_{1}^{*}=1$, then $HG_{1}^{*}/G_{1}^{*}$ satisfies the partial $\Pi$-property in $G/G_{1}^{*}$. Hence $HG_{1}^{*}/G_{1}^{*}$ has a complement, $T/G_{1}^{*}$ say, in $G/G_{1}^{*}$ by induction. It follows that $G=HG_{1}^{*}T=HT$ and $HG_{1}^{*}\cap T=G_{1}^{*}$. Since $H\cap G_{1}^{*}=1$, we have $H\cap T=1$. Hence $T$ is a complement of $H$ in $G$, as required.
(2$\Rightarrow$(1). Suppose that $H$ is complemented in $G$. We show that $H$ satisfies the $\Pi$-property in $G$ by induction on $|G|$. Let $B$ be a complement of $H$ in $G$. Then $P\cap B\unlhd G$ and $P=H\times(P\cap B)$. Since $H(P\cap B)/P\cap B$ is complemented in $G/P\cap B$, it follows that $H(P\cap B)/(P\cap B)$ satisfies the partial $\Pi$-property in $G/(P\cap B)$. Now, if $C/D$ is a $G$-chief factor below $P\cap B$, then $HD\cap C=D$ is normal in $G$. Consequently, $H$ satisfies the partial $\Pi$-property in $G$. ∎
The following lemma is a significant result proved by Zeng.
Lemma 2.13 ([13, Theorem B]).
Assume that a $p^{\prime}$-group $H$ acts faithfully on an elementary abelian $p$-group $P$. Assume that $G=P\rtimes H$ and let $d$ be a power of $p$ such that $1<d<|P|$. Then all subgroups of $P$ of order $d$ are complemented in $G$ if and only if one of the following statements holds:
(1) $G$ is supersoluble.
(2) $H$ is cyclic and $P$ is a homogeneous $\mathbb{F}_{p}[H]$-module with all its irreducible $\mathbb{F}_{p}[H]$-submodules having dimension $k>1$. Furthermore, $k$ divides $\gcd(\log_{p}d,\log_{p}|P|)$.
Lemma 2.14 ([8, Lemma 3.3(4)]).
Let $H$ be a $p^{\prime}$-group and let $V$ be a faithful and irreducible $\mathbb{F}_{p}[H]$-module. Assume that $dim_{\mathbb{F}_{p}}V$ is prime. Then $H$ is cyclic if and only if $V$ is not absolutely irreducible.
3 Proofs
We work towards the proofs of our main results.
Lemma 3.1.
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $p^{2}\leq d<|P|$. Assume that $O_{p^{\prime}}(G)=1$ and every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$. If $G$ has a minimal normal subgroup of order $d$, then $G/\mathrm{Soc}(G)$ is cyclic, $\Phi(G)=1$ and $P=\mathrm{Soc}(G)$ is a homogeneous $\mathbb{F}_{p}[G]$-module.
Proof.
By Theorem 1.2, $G$ is $p$-soluble with $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P$ is normal in $G$ and so the Schur-Zassenhaus Theorem implies that $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Since $G$ has a minimal normal subgroup of order $d$, it follows from Lemma 2.3 that every minimal normal subgroup of $G$ is elementary abelian of order $d$. Thus $N=\mathrm{Soc}(G)\leq P=O_{p}(G)$ and $N$ is a completely reducible $G$-module by Maschke’s theorem.
First we prove that $G/N$ is $p$-supersoluble. Assume not and let $T\unlhd G$ be of minimal order such that $T\leq P$ and $T/(T\cap N)\nleq Z_{{\mathcal{U}}_{p}}(G/(T\cap N))$. Then $T\cap N=V_{1}\times\cdots\times V_{s}$, $s\geq 1$, where $V_{i}$ is a minimal normal subgroup of $G$ for all $1\leq i\leq s$.
Assume that $T$ has two different maximal $G$-invariant subgroups, $U_{1}$, $U_{2}$ say. By minimality of $T$, $U_{j}/(U_{j}\cap N)\leq Z_{{\mathcal{U}}_{p}}(G/(U_{j}\cap N))$ for each $j\in\{1,2\}$. This implies that $U_{1}N/N$ and $U_{2}N/N$ are contained in $Z_{{\mathcal{U}}_{p}}(G/N)$. Consequently, $T/(T\cap N)\cong TN/N=(U_{1}N/N)(U_{2}N/N)\leq Z_{{\mathcal{U}}_{p}}(G/N)$, a contradiction. Hence $T$ has a unique maximal $G$-invariant subgroup, $U$ say. Clearly, $T>T\cap N$. Hence $U\geq T\cap N$. Since $U/(U\cap N)\leq Z_{{\mathcal{U}}_{p}}(G/(U\cap N))$ and $U\cap N=T\cap N$, we deduce that $U/(T\cap N)\leq Z_{{\mathcal{U}}_{p}}(G/(T\cap N))$. Hence $T/U\nleq Z_{{\mathcal{U}}_{p}}(G/U)$.
Write $N_{0}=T\cap N$ and denote with bars the images in $\overline{G}=G/N_{0}$. Then $\overline{U}=\overline{T}\cap Z_{{\mathcal{U}}_{p}}(\overline{G})$. We claim that the exponent of $\overline{T}$ is $p$ or $4$ (when $\overline{T}$ is not quaternion-free).
Let $\overline{D}$ be a Thompson critical subgroup of $\overline{T}$. If $\Omega(\overline{T})<\overline{T}$, then $\Omega(\overline{T})\leq Z_{{\mathcal{U}}_{p}}(\overline{G})$. By Lemma 2.5, we have that $\overline{T}\leq Z_{{\mathcal{U}}_{p}}(\overline{G})$, a contradiction. Thus $\overline{T}=\overline{D}=\Omega(\overline{D})$. If $\overline{T}$ is a non-abelian quaternion-free $2$-group, then $\overline{T}$ has a characteristic subgroup $\overline{T_{0}}$ of index $2$ by Lemma 2.7. The uniqueness of $U$ implies $T_{0}=U$. Thus $\overline{T}\leq Z_{{\mathcal{U}}_{p}}(\overline{G})$, which is impossible. Hence if $\overline{T}$ is a non-abelian $2$-group, then $\overline{T}$ is not quaternion-free. By Lemma 2.6, the exponent of $\overline{T}$ is $p$ or $4$ (when $\overline{T}$ is not quaternion-free), as claimed.
There exists a normal subgroup $\overline{Q}$ of $\overline{P}$ such that $\overline{U}<\overline{Q}<\overline{T}$ and $|\overline{Q}:\overline{U}|=p$. Let $\overline{x}\in\overline{Q}\setminus\overline{U}$. Then $\langle\overline{x}\rangle\overline{U}=\overline{Q}$ and $\langle\overline{x}\rangle$ has order $p$ or $4$. Since $N_{0}$ is elementary abelian, we have $o(x)\leq p^{2}$ or $o(x)\leq 8$.
If $o(x)>d$, then $o(x)=8$, $p=2$ and $d=4$. Since $x^{2}\in U$ and $o(x^{2})=4$, we have $\Phi(U)>1$. Let $A$ be a minimal normal subgroup of $G$ contained in $\Phi(U)$. Then $A\leq T\cap N=N_{0}$. Without loss of generality, we may assume $A=V_{1}$. Set $\widetilde{G}=G/(V_{2}\times\cdots\times V_{s})$ and denote with $\widetilde{X}$ the image of a subgroup $X$ of $G$ in $\widetilde{G}$. Clearly, $\widetilde{N_{0}}\leq\widetilde{\Phi(U)}\leq\Phi(\widetilde{U})$ and $\widetilde{U}/\widetilde{N_{0}}\leq Z_{{\mathcal{U}}_{p}}(\widetilde{G}/\widetilde{N_{0}})$. By Lemma 2.8, we have $\widetilde{U}\leq Z_{{\mathcal{U}}_{p}}(\widetilde{G})$. This implies that $V_{1}\cong\widetilde{N_{0}}$ is cyclic, a contradiction.
Suppose that $o(x)\leq d$. Let $L$ be a subroup of $Q$ of order $d$ such that $\langle x\rangle\leq L$. Then $Q=LU$. By hypothesis, $L$ satisfies the partial $\Pi$-property in $G$. Applying Lemma 2.2, $G$ has a chief series
$\varOmega_{G}:1=G^{*}_{0}<G^{*}_{1}<\cdot\cdot\cdot<G^{*}_{r-1}<G^{*}_{r}=T<\cdot\cdot\cdot<G^{*}_{n}=G$
passing through $T$ such that $|G:N_{G}(LG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. Since $U$ is the unique maximal $G$-invariant subgroup of $T$, it follows that $G_{r-1}^{*}=U$. Therefore $|G:N_{G}(LU\cap T|=|G:N_{G}(Q)|$ is a $p$-number. It follows that $Q\unlhd G$, which contradicts the fact that $U$ is the unique maximal $G$-invariant subgroup of $T$.
Consequently, $G/N$ is $p$-supersoluble. Assume that $\Phi(G)>1$. Let $B$ be a minimal normal subgroup of $G$ contained in $\Phi(G)$. Then $B\leq N=\mathrm{Soc}(G)$. Since $N$ is a completely reducible $G$-module, there exists $K\unlhd G$ such that $N=B\times K$. Observe that $N/K\leq\Phi(G)K/K\leq\Phi(G/K)$ and $(G/K)/(N/K)\cong G/N$ is $p$-supersoluble. It follows by Lemma 2.8 that $G/K$ is $p$-supersoluble, and thus $B\cong N/K$ is cyclic, a contradiction. As a consequence, $\Phi(G)=1$. Then $\mathrm{Soc}(G)=N=P$. By Lemma 2.12, every subgroup of $P$ of order $d$ is complemented in $G$. Applying Lemma 2.13, we conclude that $G/\mathrm{Soc}(G)$ is cyclic and $P$ is a homogeneous $\mathbb{F}_{p}[G]$-module. ∎
Lemma 3.2.
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$, and assume further that every cyclic subgroup of $P$ of order $4$ satisfies the partial $\Pi$-property in $G$ when $d=2$ and $P$ is not quaternion-free. If the $p$-rank of $G$ is greater than $1$, then $d\geq p^{2}|\Phi(P)|$.
Proof.
Assume the result is not true and let $G$ be a counterexample of minimal order. Then $d<p^{2}|\Phi(P)|$. We may assume that $O_{p^{\prime}}(G)=1$. By Theorem 1.2, $G$ is $p$-soluble of $p$-length at most $1$. Hence $P$ is normal in $G$ and $\mathrm{Soc}(G)\leq P$. If $d=p$, then $G$ is $p$-supersoluble by [4, Proposition 1.6], a contradiction. Hence $d\geq p^{2}$ and $\Phi(P)>1$. By Lemma 2.3(1), every minimal normal subgroup of $G$ is a $p$-group of order at most $d$. If $G$ has a minimal normal subgroup of order $d$, then $\Phi(P)=\Phi(G)=1$ by Lemma 3.1, a contradiction. Hence every minimal normal subgroup of $G$ has order at most $\frac{d}{p}$.
Let $N$ be a minimal normal subgroup of $G$ contained in $\Phi(P)$. Clearly, $\frac{d}{|N|}\geq p$ since $|N|\leq\frac{d}{p}$. As $N\leq\Phi(P)\leq\Phi(G)$ and the $p$-rank of $G$ is greater than one, we can apply [6, Theorem IV.8.6] to conclude that the $p$-rank of $G/N$ is greater than $1$. By Lemma 2.1, every subgroup of $P/N$ of order $\frac{d}{|N|}$ satisfies the partial $\Pi$-property in $G/N$. If $p\geq 3$, then $G/N$ satisfies the hypotheses of the lemma. The minimal choice of $G$ implies that $\frac{d}{|N|}\geq p^{2}|\Phi(P/N)|=p^{2}|\Phi(P)/N|$, and hence $d\geq p^{2}|\Phi(P)|$, a contradiction. This yields that $p=2$ and $|N|=\frac{d}{2}$.
Therefore every subgroup of $P/N$ of order $2$ satisfies the partial $\Pi$-property in $G/N$. By [4, Proposition 1.6], we may assume that $P/N$ is not quaternion-free. By Lemma 2.10 that every minimal normal subgroup of $G/N$ of order divisible by $2$ has order $2$. Thus every minimal normal subgroup of $G$ different from $N$ has order $2$.
If $G/N$ is $2$-supersoluble, then $G$ is $2$-supersoluble, a contradiction. Hence $G/N$ is not $2$-supersoluble. Let $M$ be a normal subgroup of $G$ of minimal order containing $N$ such that $M/N\nleq Z_{{\mathcal{U}}}(G/N)$ and $M\leq P$. Let $U/N$ be a minimal normal subgroup of $G/N$ contained in $M/N$. Then $U/N=M/N\cap Z_{{\mathcal{U}}}(G/N)$. By Lemma 2.4, $M/N$ possesses a cyclic subgroup $\langle x\rangle N/N$ of order $2$ or $4$ (when $M/N$ is not quaternion-free) such that $x\in M\setminus U$.
Denote with bars the images in $\overline{G}$. We will show that the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free). Let $\overline{D}$ be a Thompson critical subgroup of $\overline{M}$. If $\Omega(\overline{D})<\overline{M}$, then $\Omega(\overline{D})\leq\overline{U}$ by the minimal choice of $M$. Thus $\Omega(\overline{D})\leq Z_{{\mathcal{U}}}(\overline{G})$. By Lemma 2.5, we have that $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, a contradiction. Hence $\overline{M}=\overline{D}=\Omega(\overline{D})$. If $\overline{M}$ is a non-abelian quaternion-free $2$-group, then $\overline{M}$ has a characteristic subgroup $\overline{M_{0}}$ of index $2$ by Lemma 2.7. The minimal choice of $M$ implies $M_{0}=U$. Thus $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, which is impossible. Hence if $\overline{M}$ is a non-abelian $2$-group, then $\overline{M}$ is not quaternion-free. By Lemma 2.6, the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free), as desired.
Suppose that $d\geq 8$ and $N$ is the unique minimal normal subgroup of $G$. There exists a normal subgroup $\overline{Q}$ of $\overline{P}$ such that $\overline{U}<\overline{Q}<\overline{M}$ and $|\overline{Q}:\overline{U}|=2$. Let $\overline{y}\in\overline{Q}\setminus\overline{U}$. Then $|\langle\overline{y}\rangle|\leq 4$ and $\langle\overline{y}\rangle\overline{U}=\overline{Q}$. Since $N$ is elementary abelian, we have $o(y)\leq 8$. Let $L$ be a subroup of $Q$ of order $d$ such that $\langle y\rangle\leq L$. Then $Q=LU$. By hypothesis, $L$ satisfies the partial $\Pi$-property in $G$. By the uniqueness of $N$ and Lemma 2.2, $G$ has a chief series
$\varOmega_{G}:1=G^{*}_{0}<G^{*}_{1}=N<\cdot\cdot\cdot<G^{*}_{r-1}<G^{*}_{r}=M<\cdot\cdot\cdot<G^{*}_{n}=G$
passing through $M$ and $N$ such that $|G:N_{G}(LG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. The choice of $M$ yields $G_{r-1}^{*}=U$. Therefore $|G:N_{G}(LU\cap M)|=|G:N_{G}(Q)|$ is a $2$-number. It follows that $Q\unlhd G$, which contradicts the fact that $U$ is the unique maximal $G$-invariant subgroup of $M$.
Suppose that $d\geq 8$ and $N$ is not the unique minimal normal subgroup of $G$. Then $G$ has a minimal normal subgroup of order $2$, $T$ say. Then the $2$-rank of $G/T$ is greater than $1$. Note that every subgroup of $P/T$ of order $\frac{d}{2}$ satisfies the partial $\Pi$-property in $G/T$ and $\frac{d}{2}\geq 4$. The minimal choice of $G$ implies $\frac{d}{2}\geq 2^{2}|\Phi(P/T)|$. Obviously $|\Phi(P/T)|\geq\frac{|\Phi(P)|}{2}$, and it follows that $d\geq 2^{2}|\Phi(P)|$, a contradiction.
[…]
Lemma 3.3.
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$, and assume further that every cyclic subgroup of $P$ of order $4$ satisfies the partial $\Pi$-property in $G$ when $d=2$ and $P$ is not quaternion-free. Assume that $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, $P$ is elementary abelian and the $p$-rank of $G$ is greater than $1$. Then the following statements hold:
(1) Then $P$ is a homogeneous $\mathbb{F}_{p}[H]$-module with all its irreducible $\mathbb{F}_{p}[H]$-submodules having dimension $k>1$ and $k$ divides $\gcd(\log_{p}d,\log_{p}|P|)$.
(2) Every irreducible $\mathbb{F}_{p}[H]$-submodule of $P$ is not absolutely irreducible.
Proof.
It is clear that $P$ is a completely reducible $\mathbb{F}_{p}[H]$-module by Maschke’s theorem. Write $P=V_{1}\times\cdots\times V_{t}$, where $V_{i}$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P$ for all $1\leq i\leq t$. Write $\mathrm{dim}V_{i}=n_{i}$ and suppose that $n_{1}\leq n_{2}\leq\cdots\leq n_{t}$. By Lemma 2.3, we have $d\geq p^{n_{t}}$. Since $G$ is not $p$-supersoluble, it turns out that $d\geq p^{n_{t}}\geq p^{2}$ by [4, Proposition 1.6].
(1) We argue by induction on $|G|$. If $d=p^{n_{t}}$, then $n_{1}=n_{2}=\cdots=n_{t}$ by Lemma 2.3. In this case, the conclusion follows by Lemmas 2.12 and 2.13. We may assume that $d>p^{n_{t}}$ and $t>1$.
Suppose that $t=2$. By Lemma 2.1, every subgroup of $P/V_{1}$ of order $\frac{d}{|V_{1}|}$ satisfies the partial $\Pi$-property in $G/N$. Applying Lemma 2.3, we see that $|P/V_{1}|=p^{n_{2}}\leq\frac{d}{p^{n_{1}}}$. Consequently, $|P|=p^{n_{1}+n_{2}}\leq d<|P|$, a contradiction. Therefore $t\geq 3$. Note that every subgroup of $P/V_{j}$ of order $\frac{d}{|V_{j}|}$ satisfies the partial $\Pi$-property in $G/V_{j}$ and the $p$-rank of $G/V_{j}$ is equal to $n_{t}>1$ for $j=1,2$. Applying the inductive hypothesis to $G/V_{1}$ and $G/V_{2}$, we deduce that $V_{1},...,V_{t}$ are isomorphic $\mathbb{F}_{p}[H]$-modules of the same dimension $k>1$. Applying the inductive hypothesis to $G/V_{1}$, we have $k\big{|}\gcd(\log_{p}d-k,\log_{p}|P|-k)$. Hence $k\big{|}\gcd(\log_{p}d,\log_{p}|P|)$, as wanted.
(2) Let $U$ and $V$ be two isomorphic irreducible $\mathbb{F}_{p}[H]$-modules of $P$, and let $\varphi:U\mapsto V$ be an $H$-isomorphism. Suppose that $U$ is absolutely irreducible. Then $\mathrm{End}_{\mathbb{F}_{p}[H]}(U)\cong\mathbb{F}_{p}$. According to [6, Proposition B.8.2], the minimal normal subgroups of $G$ contained in $U\times V$ are exactly
$U$ and $V_{m}=\{u^{m}\varphi(u)|u\in U\}$, where $m=0,1,...,p-1$.
Let $1\not=u_{1}\in U$, and let $A$ be a subgroup of $U\times V$ of order $d$ such that $\langle u_{1},\varphi(u_{1})\rangle\subseteq A$. It is rather easy to see that $\mathrm{Core}_{G}(A)=1$. By Lemma 2.2, there exists a minimal normal subgroup $B$ of $G$ contained in $U\times V$ such that $|G:N_{G}(A\cap B)|$ is a $p$-number. Since $P$ is elementary abelian, we have $A\cap B\unlhd G$. The minimality of $B$ yields that $A\cap B=B$ or $1$. If $A\cap B=B$, then $\mathrm{Core}_{G}(A)>1$, a contradiction. Therefore $A\cap B=1$. However, $1\not=u_{1}\in A\cap U$ and $1\not=u_{1}^{m}\varphi(u_{1})\in A\cap V_{m}$ for $m=0,1,...,p-1$. This contradiction yields that $U$ is not absolutely irreducible. ∎
Lemma 3.4.
Let $G$ be a $p$-soluble group of $p$-rank greater than $1$ and let $P$ be a Sylow $p$-subgroup of $G$ with $|P|\geq p^{2}$. Suppose that every $2$-maximal subgroup of $P$ satisfies the partial $\Pi$-property in $G$. If $Q$ is a $2$-maximal subgroup of $P$, then $\Phi(P)$ is contained in $Q$.
Proof.
We argue by induction on $|G|$. By Lemma 2.1, $G/O_{p^{\prime}}(G)$ inherits the hypotheses of the lemma. Hence we may assume that $O_{p^{\prime}}(G)=1$. In view of Theorem 1.2, we have $P=O_{p}(G)=F(G)$. By Lemma 2.11, every $2$-maximal subgroup of $P$ is a partial CAP-subgroup of $G$. Applying [3, Lemmas 10 and 12], we conclude that $\Phi(P)$ is contained in $Q$. ∎
Proof of Theorem A.
Applying Theorem 1.2, we have that $G$ is $p$-soluble of $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P=O_{p}(G)$ and so $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, by the Schur-Zassenhaus Theorem.
Assume that $G$ is not $p$-supersoluble. If $|P|=p^{2}$, then $P$ is a minimal normal subgroup of $G$. Consequently, $G$ is of type (2).
Suppose that $|P|\geq p^{3}$. By Lemma 3.2, we have $\Phi(P)=1$. By Lemma 3.3(1), $P=V_{1}\times\cdot\cdot\cdot\times V_{s}$, $s\geq 2$, where $V_{i}$ is an $H$-isomorphic irreducible $\mathbb{F}_{p}[H]$-module of dimension $2$ for all $1\leq i\leq s$. Furthermore, $s\geq 2$ and $|P|\geq p^{4}$. Applying Lemma 2.14 and and Lemma 3.3(2), we have that $H$ is cyclic. Therefore, $G$ is of type (4). ∎
Proof of Theorem B.
Assume that $|P|\geq p^{2}$ and $G$ is not $p$-supersoluble. Suppose that $|P|=p^{2}$. If $G$ is $p$-soluble, then $P=O_{p}(G)$ is a minimal normal subgroup of $G$ and $G$ is of type (2).
Assume that $|P|\geq p^{3}$. If $\frac{|P|}{p^{2}}=2$, then $|P|=8$. Since $G$ is not $2$-supersoluble, we can apply Theorem 1.2 to conclude that $P\cong Q_{8}$. Hence $G$ is of type (3).
Suppose that $d=\frac{|P|}{p^{2}}>2$. By Theorem 1.2, we have that $G$ is $p$-soluble of $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P=O_{p}(G)$ and so $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, by the Schur-Zassenhaus Theorem. By Lemma 3.2, we have $\frac{|P|}{|\Phi(P)|}>\frac{d}{|\Phi(P)|}\geq p^{2}$. Applying Lemma 2.1, every $2$-maximal subgroup of $P/\Phi(P)$ satisfies the partial $\Pi$-property in $G/\Phi(P)$. By Lemma 3.3, we have $P/\Phi(P)=V_{1}\times\cdot\cdot\cdot\times V_{s}$, where $V_{i}$ is an $H$-isomorphic irreducible $\mathbb{F}_{p}[H]$-module of dimension $2$, which is not absolutely irreducible, for all $1\leq i\leq s$. It follows from 2.14 that $H$ is cyclic. Note that $|P|>d\geq p^{2}|\Phi(P)|$ and $\Phi(P)$ is contained in every $2$-maximal subgroup of $P$ by Lemma 3.4. Hence $\Phi(P)$ equals to the intersection of all $2$-maximal subgroups of $P$. Therefore $G$ is of type (4). ∎
Proof of Theorem C.
Applying Theorem 1.2, we have that $G$ is $p$-soluble of $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P=O_{p}(G)$ and so $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, by the Schur-Zassenhaus Theorem. By Lemma 3.2, we know $\frac{d}{|\Phi(P)|}\geq p^{2}$. According to Lemma 2.1, every subgroup of $P/\Phi(P)$ of order $\frac{d}{|\Phi(P)|}$ satisfies the partial $\Pi$-property in $G/\Phi(P)$. Since $P/\Phi(P)$ is elementary abelian, all subgroups of $P/\Phi(P)$ of order $\frac{d}{|\Phi(P)|}$ are complemented in $G/\Phi(P)$ by Lemma 2.12. By Lemma 3.3, $G$ satisfies Statements (1) and (2). We also have $O_{p^{\prime}}(G/\Phi(P))=1$ and so $H\Phi(P)/\Phi(P)$ acts faithfully on $P/\Phi(P)$ by [7, Theorem 6.3.2]. By Lemma 2.13, $H$ is cyclic. ∎
Acknowledgments
[…]